OpenStudy (anonymous):

Help with calculus Find the n th derivative of y, if y = (1-x^2)cosx So we're asked for an explicit formula for the n th drivative. Thanks

OpenStudy (science0229):

These kind of problem always has a pattern. I would keep differentiating it until there is a pattern.

OpenStudy (anonymous):

I can get the pattern for cosx, the problem is that x^2cosx, the derivative just gets larger end larger

OpenStudy (science0229):

\[f(x)=(1-x^2)cosx\] \[f'(x)=(cosx-x^2cosx)'=-sinx-2xcosx+x^2sinx=-2xcosx-(1-x^2)sinx\] \[f''(x)=(-2xcosx-sinx+x^2sinx)'=-2cosx+2xsinx-cosx+2xsinx+\] \[x^2cosx\]

OpenStudy (science0229):

The second derivative simplifies farther to \[4xsinx-(3-x^2)cosx\]

OpenStudy (science0229):

The third derivative skipping the calculation, is \[6xcosx+(7-x^2)sinx\]

OpenStudy (science0229):

The fourth derivative is \[-8xsinx+(13-x^2)cosx\]

OpenStudy (science0229):

The fifth derivative is \[-10xcosx-(21-x^2)sinx\]

OpenStudy (science0229):

See a pattern?

OpenStudy (anonymous):

Maybe f^n(x) = 2nxsin(x+n*pi/2)+(somenumber-x^2)cos(x+n*pi/2)

OpenStudy (anonymous):

but what is that number

OpenStudy (science0229):

that "somenumber" are 1,3,7,13,21...etc 1+2=3 3+4=7 7+6=13 13+8=21 ...

OpenStudy (anonymous):

ok, i got it, thanks

OpenStudy (science0229):

Welcome:)